OpenStudy (anonymous):

change the cartesian integral into polar integral: integral -a to a, integral -(a^2-x^2)^1/2 to (a^2-x^2)^1/2 dy dx

OpenStudy (anonymous):

all you have to do is substitute x=rCos(theta) and y=rSin(theta) and change your integral to r*dr*dtheta

OpenStudy (anonymous):

tried that

OpenStudy (anonymous):

i got 1/2 a^2- 1/2a^2

OpenStudy (anonymous):

r=+-a

OpenStudy (anonymous):

what is your function? y=(a^2-x^2)^1/2 or is this equal to a constant?

OpenStudy (anonymous):

y^2=a^2-x^2 y^2+x^2=a&2 R^2=a^2 r=+/- a

OpenStudy (anonymous):

so my first integral becomes -a to a of r dr? correct?

OpenStudy (anonymous):

so my first integral becomes -a to a of r dr? correct?

OpenStudy (anonymous):

so this is just a circle? with radius a? if so then you have int(0,2*pi)int(0,a)r*dr*dtheta

OpenStudy (anonymous):

so this is just a circle? with radius a? if so then you have int(0,2*pi)int(0,a)r*dr*dtheta

OpenStudy (anonymous):

OMG!! r=a is a circle from 0 to a, you are right! I was thinking it was from -a to a which is a diameter not a radius, thank you!!!

OpenStudy (anonymous):

not a problem, anytime